### 14 Dec 2023 v3

@@ Abstract
This article intends to systematically identify and comparatively analyze state-of-the-art supply chain (SC) forecasting strategies and technologies. A novel framework has been proposed incorporating Big Data Analytics in SC Management (problem identification, data sources, exploratory data analysis, machine-learning model training, hyperparameter tuning, performance evaluation, and optimization), forecasting effects on human-workforce, inventory, and overall SC. Initially, the need to collect data according to SC strategy and how to collect them has been discussed. The article discusses the need for different types of forecasting according to the period or SC objective. The SC KPIs and the error-measurement systems have been recommended to optimize the top-performing model. The adverse effects of phantom inventory on forecasting and the dependence of managerial decisions on the SC KPIs for determining model performance parameters and improving operations management, transparency, and planning efficiency have been illustrated. The cyclic connection within the framework introduces preprocessing optimization based on the post-process KPIs, optimizing the overall control process (inventory management, workforce determination, cost, production and capacity planning). The contribution of this research lies in the standard SC process framework proposal, recommended forecasting data analysis, forecasting effects on SC performance, machine learning algorithms optimization followed, and in shedding light on future research.
@@ 1 Introduction
The supply chain (SC) has evolved sufficiently over the past years to discover new methods and techniques for solving SCM problems. The SC can develop its configuration based on its control, coordination, and management [89]. The advent of Big Data (BD) brings one such change. Like other fields, BD can be utilized to improve decision-making reprocesses and alter business models through multiple resources, tools, and applications [136]. Therefore, SC and BD usages are connected to help one another. Although the concepts for SCM are already well-developed, it is possible to improve further; recent researches on enhancing efficiency through collaboration [117], usage of RFID and intelligent goods [62] are two examples of such innovations that improved SCM processes. Newer technologies are further enabling the discovery of innovative strategies for solving SC problems. Big Data Analytics (BDA) is one such disruptive innovation. Although BD has been present for a long time, the approaches to making sense of BD are comparatively new, and such systems have not been wholly integrated into other branches of knowledge [148]. We identified the absence of data usage and relevant processes in SC directly as a major problem that needs to be addressed.
BD has similarly grown popular over the years. After academic and technical publications first mentioned such technological developments, it has drawn the attention of various people, including literary scholars, corporate leaders, and government officials [59]. The most recognizable feature of BD is probably its size or the amount of data stored. Beyond this feature of volume, BD is recognizable by the greater variety in data types, high data collection and analysis speed or velocity, veracity dealing necessity, random variabilities in data, and the growth in value as more data are analyzed [42]. Simply having access to BD is not helpful; data analytics is a must to create value or gather information out of the enormous collection of data. Where analytical methods are applied to BD, it is called BDA. While there are vast applications of BDA, its role in improving the SC process is notable.
@@ 1.1 Research gaps
We found potential research gaps combining the preprocessing for machine learning (ML) forecasting, control process (SC processes where BDA is helpful), and post-process (for the evaluation of the forecasting model). Although there is fragmented research on these particular topics (“SC forecasting model performance,” “the application of BDA on SC,” or “ML forecasting techniques with BDA implementation,” or “BD driven SC performance evaluation”), the necessity to form a cyclic connection among these three processes led us to the development of this article.
@@ 1.2 Research objectives
The novelty of this paper lies in several key aspects:
@@
By incorporating these novel aspects, the paper contributes to the existing body of knowledge in the field of BDA-SCM Framework for Forecasting, offering new insights, methodologies, and recommendations for future research and practical implementation.
@@ 2.1 Planning the review
In this article, the BDA-SCM cyclic framework was developed initially incorporating pre-process, control process, and post-process phases. Each phase was illustrated utilizing the most relevant selected works of literature. For forecasting purposes, pre-process recommendations include a step-by-step approach to forecasting and BDA best practices to facilitate comprehensive demand forecasting considering state-of-the-art technologies and relevant research. In the control process, how SC factors and forecasting affect workforce efficiency have been discussed. The post-process portion of the managerial decision-making process explains how managers use the KPI and optimization of the forecasting model to choose the appropriate metrics and insights.
@@ 2.2.2 Selection strategy
To ensure that the selected papers were empirically sound and conceptually relevant to BDA-SCM-related research advances, the relevance of each publication was assessed. Articles were considered more relevant if the search terms appeared in the title, abstract, keywords, and throughout the text. The identified papers were critically analyzed, particularly in terms of the relevant sections that mentioned BDA-SCM. This approach drew from relevant views on SCM-forecasting challenges and BDA techniques and helped to achieve the research review goals. The remaining articles were then assessed to verify that they provided the necessary research perspective and empirical data to meet the review’s objectives. Finally, to ensure that the selected articles aligned with the review goals, we conducted a rigorous alignment process, comparing the articles to the research review objectives. Only articles that met all of the selection criteria were included in the final review.
@@ 3 BDA-SCM framework
Figure 2 mainly consists of the use and cyclic flow of data in SC. It only includes the SC parts where BDA may be involved. Figure 3 complements Figure 2 by mentioning the methods for cleaning, exploring, and analyzing data properly. It includes feature engineering techniques to select only the most relevant and unique features from which ML algorithms can learn efficiently. Finally, Figure 4 is a proposed method for data splitting, model training, hyperparameter optimization, cross-validation, testing, and evaluating errors to perfect the forecasting methods mentioned in Figure 2.
@@ 4.1 Identifying business problems
At the outset, the type of data that needs to be collected, stored, analyzed, and interpreted is selected based on SC strategies. [143] categorized SC strategies based on risk and impact, such as robustness for low-impact high-risk, agility for low-risk high-impact, rigidity for low-impact low-risk, and resilience for high-impact high-risk decisions. The strategies can be adopted by varying levels of responsiveness and efficiency. Responsiveness has been a critical factor in gaining a competitive advantage, and it depends on the deviations in demands and a company’s capability to respond to such deviations. An increase in responsiveness decreases efficiency, and vice versa [93]. The responsiveness level affects product volume, order fulfillment rate, workforce, manufacturing capacity, warehouse capacity, transportation carriers, product mix, supplier’s product mix, inbound and outbound logistics, etc. [128]. Therefore, whether data needs to be collected should be decided based on responsiveness, as different sets of data by allocating are required to boost the SC efficiency and responsiveness. Furthermore, the frequency of data analyses would also depend on the responsiveness level of the company. In short, the factors that may dictate the sort of forecasting required for a business include the context of forecasting, the types of data available, the required level of accuracy, the length of the forecasting period, the time available for each forecast, and the value addition made through the forecast [20].
@@ 4.3.1 Duplicates removal
It is problematic to waste space and runtime with duplicate rows. Duplicate rows create incoherence, and the ML model fails to learn new information. Because of input mistakes, changes in some feature values (e.g., the identifier value) may generate duplicate rows that will be deemed distinct by the machine. Using data preprocessing libraries in Python and R, it is easier to drop the duplicates or substitute them with relevant values. Nevertheless, the main challenge is the identification of factors on which the duplicates should be removed. Of the number of methods invented to remove duplicates, we review the following:
Bayesian: The Fellegi-Sunter-algorithm is the most commonly used model in probabilistic approaches because of its Bayesian nature [37]. The Bayes Decision Rule is a common approach [33]. A Bayesian inference difficulty may develop when the probability density of a unique row differs from a duplicate record and the functions are known. Neural network (NN) algorithms are more accurate without the Fellegi-Sunter-algorithm if the data are adequately described or labeled [144].
@@ 4.3.2 Dealing with categorical features
One-Hot, Ordinal, Helmert, Polynomial, and Binary encoders are outperformed with a 95% accuracy by Sum, and Backward Difference encoders are preferred for prediction jobs [102]. [87] presented a generic Information-based encoder that transforms mixed-type features into numeric ones while maintaining the dataset’s original dimension, with better accuracy than One-Hot and Feature-Hashing. [47] demonstrated that Ordinal-encoder (straightforward and convenient to execute but incorporates a sequence of features) outperformed Hashing (introduces limited features and moderately ignores the feature sequence); One-hot-encoding generates a massive number of features and forces the use of a very simplified regression analysis. To train residual features from time categorical variables derived from variable time stamps, DeepGB neural network with embedding layers may be used, which are necessary to learn multiple time series at once to encode categorical features in a lower dimension or by embedding their IDs and retrieve helpful information [73].
@@ 4.3.4 Data transformation
Data transformation can create new features, also known as changing features, where mathematical formulae derived from business models or pure mathematical formulae are used to integrate the raw input features. Linear, quadratic, polynomial, non-polynomial, rank and Box-Cox transformations are a few to name among different existing transformation techniques.
@@
Nonparametric approaches are not recommended to be introduced into traditional statistics courses using rank transformation because it inhibits how widely the nonparametric technique may be used, which is unnecessary. Another misperception is that the nonparametric technique is utilized chiefly for hypothesis testing. This entirely obscures the superior theoretical and conceptual flexibility of many nonparametric methods.
@@ 4.3.5 Filtering extreme outliers
Outlier treatment:
Trimming: Outliers from the dataset are removed, and it’s not recommended.
Capping: In capping, if the values are greater or less than an established threshold, those are deemed outliers, and the outlier numbers in the dataset determine the capping number.
Treat outliers as MVs: Outliers may be handled as the MVs are treated.
Outlier Removal Clustering (ORC) eliminates outliers in each loop, a modification of K-Means Clustering. ORC efficiently removes outliers from clusters. The parameters should be adjusted appropriately as the model precision changes depending on the dataset. ORC guarantees that the centroid computation is not biased by locations distant from the k-clusters.
@@ 4.3.6 Dealing with missing values (MVs)
Maximum likelihood imputation methods: Assume for $n$ independent rows $(i=1,…,n)$, there are $k$ variables $(y_{i}1,y_{i}2,…,y_{i}k)$ with no missing data. The maximum likelihood function is:
@@ 4.3.7 Binning
In this method, a continuous variable is converted into a group of intervals. Each interval can then be treated as a ‘bin,’ with the option of enforcing an order dependent on the data’s subsequent processing. While smoothing, each bin’s min and max values are calculated as bin borders. Then, for each value, the nearest border value is substituted. Typically, the smoothing effect increases with bin width. If the bin widths are identical, binning may be used as a discretization method by substituting mean or median for bin value. It is possible to create hierarchical ideas by iterating over this procedure indefinitely. It’s unsupervised since class labels are not used, and bin numbers are specified by the user.
4.3.8 Deep-Learning (DL) based FE

### 07 Feb 2024 v5

@@ Abstract
This article systematically identifies and comparatively analyzes state-of-the-art supply chain (SC) forecasting strategies and technologies within a specific timeframe, encompassing a comprehensive review of 152 papers spanning from 1969 to 2023. A novel framework has been proposed incorporating Big Data Analytics in SC Management (problem identification, data sources, exploratory data analysis, machine-learning model training, hyperparameter tuning, performance evaluation, and optimization), forecasting effects on human workforce, inventory, and overall SC. Initially, the need to collect data according to SC strategy and how to collect them has been discussed. The article discusses the need for different types of forecasting according to the period or SC objective. The SC KPIs and the error-measurement systems have been recommended to optimize the top-performing model. The adverse effects of phantom inventory on forecasting and the dependence of managerial decisions on the SC KPIs for determining model performance parameters and improving operations management, transparency, and planning efficiency have been illustrated. The cyclic connection within the framework introduces preprocessing optimization based on the post-process KPIs, optimizing the overall control process (inventory management, workforce determination, cost, production & capacity planning). The contribution of this research lies in the standard SC process framework proposal, recommended forecasting data analysis, forecasting effects on SC performance, machine learning algorithms optimization followed, and in shedding light on future research.
@@ 1 Introduction
In the dynamic landscape of supply chain management (SCM), the relentless pursuit of efficiency and adaptability has driven a continuous evolution in forecasting strategies and technologies. This article embarks on a systematic exploration, aiming to identify and analyze the state-of-the-art in supply chain (SC) forecasting, ultimately proposing a novel framework that integrates the power of big data analytics (BDA) into SCM. The increasing complexity and interconnectedness of global SCs have underscored the need for sophisticated forecasting strategies. Traditional approaches are being reevaluated in the wake of technological advancements, leading to a paradigm shift in how we perceive and optimize SC forecasting. The integration of BDA emerges as a transformative force, promising enhanced predictive capabilities and a holistic framework that spans problem identification, data sourcing, exploratory data analysis, machine learning (ML) model training, hyperparameter tuning, performance evaluation, and optimization.
The SC has evolved sufficiently over the past years to discover new methods and techniques for solving SCM problems. The SC can develop its configuration based on its control, coordination, and management [89]. The advent of big data (BD) brings one such change. Like other fields, BD can be utilized to improve decision-making reprocesses and alter business models through multiple resources, tools, and applications [136]. Therefore, SC and BD usages are connected to help one another. Although the concepts for SCM are already well-developed, it is possible to improve further; recent research on enhancing efficiency through collaboration [117], usage of RFID and intelligent goods [62] are two examples of such innovations that improved SCM processes. Newer technologies are further enabling the discovery of innovative strategies for solving SC problems. BDA is one such disruptive innovation. Although BD has been present for a long time, the approaches to making sense of BD are comparatively new, and such systems have not been wholly integrated into other branches of knowledge [148]. We identified the absence of data usage and relevant processes in SC directly as a major problem that needs to be addressed.
BD has similarly grown popular over the years. After academic and technical publications first mentioned such technological developments, it drew the attention of various people, including literary scholars, corporate leaders, and government officials [59]. The most recognizable feature of BD is probably its size or the amount of data stored. The distinctive features of greater data variety, high velocity in collection and analysis, the necessity to navigate veracity challenges, and the inherent value growth with increasing data analysis [42] set the stage for a new era in SCM. Simply having access to BD is not helpful; data analytics is a must to create value or gather information out of the enormous collection of data. Where analytical methods are applied to BD, it is called BDA. While BDA has vast applications, its role in improving the SC process is notable.
The motivation behind this research stems from the recognition that the traditional forecasting paradigms may not be equipped to address the intricacies of modern SC dynamics. Phantom inventory, varying time horizons, and diverse SCM objectives necessitate a more adaptive and data-driven approach. By introducing a comprehensive framework, this article seeks to bridge the gap between traditional forecasting methods and the demands of contemporary SC environments.
@@ 1.1 Research gaps
We found potential research gaps combining the preprocessing for ML forecasting, control process (SC processes where BDA is helpful), and post-process (for the evaluation of the forecasting model). Although there is fragmented research on these particular topics (“SC forecasting model performance,” “the application of BDA on SC,” or “ML forecasting techniques with BDA implementation,” or “BD driven SC performance evaluation”), the necessity to form a cyclic connection among these three processes led us to the development of this article. This paper identifies a critical gap in current SCM practices — the underutilization of data and relevant processes - and positions BDA as a powerful solution. The motivation behind this study lies in addressing the pivotal challenge of harnessing the full potential of BD within SCM, recognizing it not just as a technological evolution but as a strategic imperative for future competitiveness.
@@ 1.2 Research objectives
The scope of this research extends beyond the technical intricacies of forecasting algorithms. It delves into the broader implications of forecasting on the human workforce, inventory management, and the overall performance of the SC. By addressing the adverse effects of phantom inventory and emphasizing the dependency of managerial decisions on SC key performance indicators (KPIs), this research contributes to the overarching goal of improving operations management, transparency, and planning efficiency. The novelty of this paper lies in several key aspects:
@@
The forthcoming sections will comprehensively explore the “Research procedures,” and the proposed cyclic connection embedded within the framework will be elucidated in “BDA-SCM framework”. Moving forward, “Pre-process” section addresses critical components, including the imperative need for strategic data collection aligned with SC objectives, methodologies for data preprocessing, feature engineering (FE), exploratory data analysis, and the classification of forecasting types based on distinct time horizons. The “Control-process” section delves into optimizing preprocessing methodologies. This optimization, rooted in post-process KPIs, aims to elevate overall control processes, encompassing inventory management, workforce determination, cost optimization, and production and capacity planning. The discussion on SC KPIs and error-measurement systems for model optimization unfolds in the “Post-process” section. This section aims to provide insights into refining forecasting models for superior performance. In the “Challenges” section, attention is directed toward acknowledging and addressing technological obstacles encountered during the extensive review of pertinent articles. This section highlights and discusses the challenges inherent in navigating the technological landscape within the scope of this research. In the “Practical Implications” section, we delve into actionable insights for SC practitioners, detailing the implementation of the proposed BDA-SCM framework in real-world scenarios and outlining the substantial benefits they can expect. By incorporating these novel aspects, the paper contributes to the existing body of knowledge in the field of BDA-SCM framework for forecasting, offering new insights, methodologies, and recommendations for future research and practical implementation.
@@ 2.1 Planning the review
In this article, the BDA-SCM cyclic framework was initially developed to incorporate pre-process, control process, and post-process phases. Each phase was illustrated utilizing the most relevant selected works of literature. For forecasting purposes, pre-process recommendations include a step-by-step approach to forecasting and BDA best practices to facilitate comprehensive demand forecasting considering state-of-the-art technologies and relevant research. In the control process, how SC factors and forecasting affect workforce efficiency have been discussed. The post-process portion of the managerial decision-making process explains how managers use the KPI and optimization of the forecasting model to choose the appropriate metrics and insights.
@@ 2.2.2 Selection strategy
The relevance of each publication was assessed to ensure that the selected papers were empirically sound and conceptually relevant to BDA-SCM-related research advances. Articles were considered more relevant if the search terms appeared in the title, abstract, keywords, and throughout the text. The identified papers were critically analyzed, particularly regarding the relevant sections that mentioned BDA-SCM. This approach drew from relevant views on SCM-forecasting challenges and BDA techniques and helped to achieve the research review goals. The remaining articles were then assessed to verify that they provided the necessary research perspective and empirical data to meet the review’s objectives. Finally, to ensure that the selected articles aligned with the review goals, we conducted a rigorous alignment process, comparing the articles to the research review objectives. Only articles that met all of the selection criteria were included in the final review.
@@ 3 BDA-SCM framework
The proposed BDA-SCM framework, depicted in Figures 2, 3, and 4, establishes a cyclic connection that facilitates continuous improvement in SC forecasting. This cyclic process seamlessly integrates three essential stages: Pre-process, Control-process, and Post-process, fostering a dynamic relationship that optimizes SC operations iteratively. Figure 2 mainly consists of the use and cyclic flow of data in SC. It only includes the SC parts where BDA may be involved. Figure 3 complements Figure 2 by mentioning the methods for cleaning, exploring, and analyzing data properly. It includes FE techniques to select only the most relevant and unique features from which ML algorithms can learn efficiently. Finally, Figure 4 is a proposed method for data splitting, model training, hyperparameter optimization, cross-validation, testing, and evaluating errors to perfect the forecasting methods mentioned in Figure 2.
In the Pre-process stage, the focus is on ensuring accurate and relevant data aligned with SC objectives. The cyclic nature of this stage involves a continuous feedback loop. For example, after training an initial ML forecasting model, the performance is evaluated using real-time data. Any discrepancies or deviations from expected outcomes trigger a revisit to the Pre-process stage. This might involve reassessing data collection methods, exploring new data sources, or refining the preprocessing steps to enhance the quality of input data. The Control-process stage benefits from the cyclic connection, encompassing decision-making areas like production planning, workforce determination, and inventory management. Suppose a decision made based on forecasted data results in suboptimal outcomes. In that case, this feedback loops back to the Pre-process stage. The system may reevaluate the forecasting model’s inputs, incorporating real-time data to enhance decision-making accuracy in subsequent cycles. In the Post-process stage, the cyclic connection enables continuous performance improvement. After the initial model predictions, performance metrics are analyzed, and any deviations from expected results trigger a reevaluation of the forecasting model. This feedback loop, integrated into the Post-process stage, ensures that the model evolves over time, adapting to changing SC dynamics and improving its predictive capabilities.
Consider a scenario in demand forecasting where the initial ML model predicts a surge in demand for a particular product. If the actual demand deviates from the forecast, the cyclic connection triggers a reassessment in the Pre-process stage. Analysts may explore new data sources, refine data preprocessing methods, or adjust FE techniques to capture changing demand patterns more accurately. In inventory management, the Control-process stage involves decisions on stock levels based on forecasted demand. If the actual inventory levels deviate significantly from the forecast, the cyclic connection prompts a revisit to the Pre-process stage. This may involve refining the preprocessing of inventory data, incorporating real-time data on SC disruptions, or adjusting the forecasting model to enhance inventory optimization. In production planning, the decision-making process relies on accurate product demand forecasts. If the actual production output falls short or exceeds the forecasted demand, the cyclic connection triggers a reassessment in the Pre-process stage. This may involve refining data collection methods, exploring new features relevant to production efficiency, or adjusting the forecasting model to better align with dynamic production needs.
@@ 4.1 Identifying business problems
At the outset, the type of data that needs to be collected, stored, analyzed, and interpreted is selected based on SC strategies. [143] categorized SC strategies based on risk and impact, such as robustness for low-impact high-risk, agility for low-risk high-impact, rigidity for low-impact low-risk, and resilience for high-impact high-risk decisions. Varying levels of responsiveness and efficiency can adopt the strategies. Responsiveness has been a critical factor in gaining a competitive advantage, and it depends on the deviations in demands and a company’s capability to respond to such deviations. An increase in responsiveness decreases efficiency, and vice versa [93]. The responsiveness level affects product volume, order fulfillment rate, workforce, manufacturing capacity, warehouse capacity, transportation carriers, product mix, supplier’s product mix, inbound and outbound logistics, etc. [128]. Therefore, whether data needs to be collected should be decided based on responsiveness, as different data sets are required to boost the SC efficiency and responsiveness by allocating them. Furthermore, the frequency of data analyses would also depend on the responsiveness level of the company. In short, the factors that may dictate the sort of forecasting required for a business include the context of forecasting, the types of data available, the required level of accuracy, the length of the forecasting period, the time available for each forecast, and the value addition made through the forecast [20].
@@ 4.3.1 Duplicates removal
It is problematic to waste space and runtime with duplicate rows. Duplicate rows create incoherence, and the ML model fails to learn new information. Because of input mistakes, changes in some feature values (e.g., the identifier value) may generate duplicate rows that will be deemed distinct by the machine. It is easier to drop the duplicates or substitute them with relevant values using data preprocessing libraries in Python and R languages. Nevertheless, the main challenge is identifying factors on which the duplicates should be removed. Of the number of methods invented to remove duplicates, we review the following:
Bayesian: The Fellegi-Sunter-algorithm is the most commonly used model in probabilistic approaches because of its Bayesian nature [37]. The Bayes Decision Rule is a common approach [33]. A Bayesian inference difficulty may develop when the probability density of a unique row differs from a duplicate record, and the functions are known. Neural network (NN) algorithms are more accurate without the Fellegi-Sunter-algorithm if the data are adequately described or labeled [144].
@@ 4.3.2 Dealing with categorical features
One-hot, ordinal, Helmert, polynomial, and binary encoders are outperformed with a 95% accuracy by Sum, and Backward Difference encoders are preferred for prediction jobs [102]. [87] presented a generic Information-based encoder that transforms mixed-type features into numeric ones while maintaining the dataset’s original dimension, with better accuracy than One-Hot and Feature-Hashing. [47] demonstrated that Ordinal-encoder (straightforward and convenient to execute but incorporates a sequence of features) outperformed Hashing (introduces limited features and moderately ignores the feature sequence); One-hot-encoding generates a massive number of features and forces the use of a very simplified regression analysis. To train residual features from time categorical variables derived from variable time stamps, a DeepGB neural network with embedding layers may be used, which are necessary to learn multiple time series at once to encode categorical features in a lower dimension or by embedding their IDs and retrieve helpful information [73].
@@ 4.3.4 Data transformation
Data transformation can create new features, also known as changing features, where mathematical formulae derived from business models or pure mathematical formulae are used to integrate the raw input features. Linear, quadratic, polynomial, non-polynomial, rank, and Box-Cox transformations are a few of the different existing transformation techniques.
@@
Nonparametric approaches using rank transformation are not recommended to be introduced into traditional statistics courses because it inhibits how widely the nonparametric technique may be used, which is unnecessary. Another misperception is that the nonparametric technique is utilized chiefly for hypothesis testing. This entirely obscures the superior theoretical and conceptual flexibility of many nonparametric methods.
@@ 4.3.5 Filtering extreme outliers
Outlier Treatment: Various techniques can be employed to address outliers within a dataset. Trimming, the first method, involves the removal of outliers, but it is generally not recommended due to potential information loss. As the second approach, capping identifies outliers based on a predefined threshold, either greater or less than the established limit. The number of outliers in the dataset influences the determination of this capping threshold. Alternatively, outliers may be treated similarly to missing values (MVs). Lastly, Outlier Removal Clustering (ORC), a modification of K-Means Clustering, eliminates outliers in iterative loops. ORC effectively removes outliers from clusters, and careful parameter adjustments are essential as the dataset influences model precision. Importantly, ORC ensures that the computation of centroids remains unbiased, particularly when dealing with distant locations from the k-clusters.
@@ 4.3.6 Dealing with missing values (MVs)
Maximum likelihood imputation methods: Assume for $n$ independent rows $(i=1,\ldots,n)$, there are $k$ variables $(y_{i}1,y_{i}2,\ldots,y_{i}k)$ with no missing data. The maximum likelihood function is:
@@ 4.3.7 Binning
In this method, a continuous variable is converted into a group of intervals. Each interval can then be treated as a ‘bin,’ with the option of enforcing an order dependent on the data’s subsequent processing. While smoothing, each bin’s min and max values are calculated as bin borders. Then, for each value, the nearest border value is substituted. Typically, the smoothing effect increases with bin width. If the bin widths are identical, binning may be used as a discretization method by substituting mean or median for bin value. It is possible to create hierarchical ideas by iterating over this procedure indefinitely. It’s unsupervised since class labels are not used, and the user specifies bin numbers.
4.3.8 Deep learning (DL) based FE